OpenStudy (saifoo.khan):

Completing square form, with steps: -0.014x^2+ 0.56x+14.4

OpenStudy (anonymous):

i hate decimals

OpenStudy (anonymous):

make it fraction

OpenStudy (anonymous):

Tofu, 9 more !

hero (hero):

divide each term by -0.014 first

OpenStudy (anonymous):

tofu, i love it really

OpenStudy (anonymous):

XD

OpenStudy (mimi_x3):

LOL, Tofu, Saifi sounds way~ BETTER :P

OpenStudy (saifoo.khan):

lol @ #s No mini it dosnt.

hero (hero):

Saifoo sucks I mean... Saifoo, this problem sucks

OpenStudy (mimi_x3):

Ofc it does, you don't deserve to be called a delicious food :P

OpenStudy (anonymous):

\[-.014(x^2-40x)+14.4\] is step one

OpenStudy (anonymous):

yea it really does suck doesn't it!

OpenStudy (saifoo.khan):

thenn....??? @sat

OpenStudy (anonymous):

i still like the name seafood better

OpenStudy (anonymous):

\[-.014(x-20)^2+14.4-.014\times 400\]

myininaya (myininaya):

i don't think that is right

OpenStudy (anonymous):

really? i may have made a mistake but i get \[-.014(x-20)^2+8.8\]

myininaya (myininaya):

let me try

OpenStudy (saifoo.khan):

it's not 8.8

OpenStudy (anonymous):

oh cramp i subtracted when i should have added sorry

OpenStudy (anonymous):

\[-.014(x-20)^2+14.4+.014\times 400\] \[-.014(x-20)^2+20\]

OpenStudy (saifoo.khan):

from where you get: 14.4+.014×400

OpenStudy (anonymous):

i forget that the first term was negative. so you were actually subtracting \[-.014\times 400\] so you have to add it back. my mistake

OpenStudy (anonymous):

oh when you "complete the square" i changed \[x^2-40\] into \[(x-20)^2=x^2-40+400\] i actually added 400

OpenStudy (anonymous):

but i didn't really add 400 because it was all multiplied by -.014

myininaya (myininaya):

\[-0.014x^2+0.56x+14.4\] \[=-0.014(x^2+\frac{0.56}{-0.014}x)+14.4\] \[=-0.014(x^2+\frac{0.56}{-0.014}x+(\frac{0.56}{2 \cdot (-0.014)})^2)+0.014 \cdot (\frac{0.56}{2 \cdot (-0.014)})^2\] \[=-0.014(x+\frac{0.56}{2 \cdot (-0.014)})^2+0.014 \cdot (\frac{0.56}{2 \cdot (-0.014)})^2\] =-0.014(x-20)^2+5.6 is that right?

OpenStudy (anonymous):

no i am right sorry

hero (hero):

What arrogance

OpenStudy (anonymous):

thank you!

myininaya (myininaya):

oops i forgot about the 14.4

hero (hero):

It's like saying "no, sorry, I'm God"

OpenStudy (anonymous):

because myinninay know i am not really sorry, i am just gloating

myininaya (myininaya):

satellite is a math god

OpenStudy (anonymous):

lol

myininaya (myininaya):

stupid decimals too much decimals

hero (hero):

No, it doesn't work that way. Either you're God, or you're not.

hero (hero):

Either we're all God, or nobody is

myininaya (myininaya):

satellite is a math god

myininaya (myininaya):

why it repeat what i said that was weird

myininaya (myininaya):

satellite is a math god

myininaya (myininaya):

it did it again :(

OpenStudy (anonymous):

@saifoo the 14.4 was in the problem, and the rest was reconciling from completing the square. that is why the last term was \[14.4+0.14\times 400\]

OpenStudy (anonymous):

satellite using his divine power

hero (hero):

Myininaya, don't do that again else God will spank you

myininaya (myininaya):

its putting that with every reply i'm not doing it

hero (hero):

It didn't put it this time. This shows that you truly fear God

myininaya (myininaya):

ok it looks like it is over satellite and i have been competing and it looks like satellite>myininaya

myininaya (myininaya):

i still didn't figure out that one with the inflection tangent

hero (hero):

I didn't think it was worth competing over. I mean, you have to consider who posted the problem. You have to ask yourself, "Is Saifoo worth competing over?".

myininaya (myininaya):

no we aren't competing over saifoo

hero (hero):

So you're not competing over Saifoo, but over his problems instead.